# Template-Based Visual Program Distillation

### v2

@@ Abstract
For users with limited computational resources, visual programming or prompting large language models (LLMs) to generate executable code for visual tasks, like visual question answering (VQA), remains largely inaccessible. Even with techniques such as distillation, adapting visual programming to smaller models or specific datasets is still quite challenging due to high annotation costs. We propose a low-cost visual program distillation method that can be used for models with fewer than 1 billion parameters and requires no human-generated program annotations. We achieve this through synthetic data augmentation based on decoupling programs into higher-level skills, called templates, and their corresponding arguments. Experimental results show that, with a relatively small amount of question/answer data, small language models can generate high-quality visual programs with the added benefit of much faster inference.
Template-Based Visual Program Distillation
@@ 1 Introduction
Visual programming Gupta and Kembhavi (2023); Subramanian et al. (2023); Surís et al. (2023) refers to generating programs that invoke visual models to solve tasks such as answering questions about images, typically by prompting a very large language model (LLM) like GPT Achiam et al. (2023) or Llama Touvron et al. (2023). Despite success in zero-shot settings, visual programming remains largely inaccessible due to high program generation costs, which come from the need for expensive hardware to run LLMs or using expensive proprietary APIs, as well as long inference time and high data costs for both human-generated program annotations and question/answer annotations. Previous efforts Khan et al. (2024) around generating visual programs on smaller LLMs have made some progress in addressing these concerns, but training and inference costs remain substantial.
Our goal is to create a visual programming system with two key characteristics:
@@
    Small Program Generator Programs should be generated by models with $<1$ billion parameters, enabling fast inference and use on a wide variety of hardware.
@@
    Limited Human Annotations Minimize the number of required annotations, especially human-generated programs.
To achieve this, our approach builds on the key observation that many visual programs share a common high-level structure – a unique aspect of visual programming. For example, the programs ‘Count the red chairs’ and ‘Count the green apples’ follow the same pattern but differ in their inputs. Inspired by this insight and recent advances in LLM distillation Hsieh et al. (2023), we propose a low-cost visual program distillation method. Our approach requires no human-generated example programs and uses only a small fraction of question/answer pairs (at most 1000). A teacher model leverages auto-context generation, where each generated program that produces the correct answer is added to the set of in-context examples. The resulting annotated programs form a dataset used to train a small language model (SLM) to generate visual programs efficiently.
@@
Distill Step-by-Step Hsieh et al. (2023) Text Gen. LLM Distillation 220M-770M 12.5% Q/A, Reasoning
@@
Template-Based Aug. (Ours) Program Gen. LLM Distillation 770M 0.1% ($\approx$1K) Visual Prog.
@@
During training, program annotations are synthetically generated through template-based augmentation. This strategy decouples high-level skills or plans, called templates, from question-specific concepts or arguments. For example, the template for ‘Count the red chairs’ is ‘find(arg1), verify property(arg2), count’ and the corresponding arguments are ‘red’ (arg1) and ‘chair’ (arg2). New question/program pairs are created by replacing these arguments in the question and program. By replacing ‘red’ with ‘green’ and ‘chairs’ with ‘apples,’ we generate both a new question and a corresponding program. Compared to directly training on the program annotations, training with template-based augmentation offers three key benefits: 1) better performance when less data is available, 2) increase in student/teacher prediction agreement, and 3) higher-quality generated programs. Our method uses a low-cost teacher model (e.g. GPT-Mini-4o) and costs only around \$1 per dataset. In addition, the distilled models achieve much faster inference speed than existing LLMs.
To validate our approach, we evaluate on frequently used visual question answering (VQA) datasets and compare with using human-generated annotations as well as non-augmented distillation. Moreover, we address an overlooked evaluation aspect by performing several types of evaluation, including human program verification, due to the noisiness of commonly used metrics.
All together, our results show that it is possible with few question/answer pairs to distill the visual program ability of LLMs into a much smaller model that does not require in-context examples.
@@ 2 Related Works
A long line of work investigates generating and executing programs to perform visual tasks. Early approaches (Andreas et al., 2015, 2016; Hu et al., 2017; Johnson et al., 2017) generate programs and execute the programs with learned end-to-end neural modules. Based on the impressive code generation capabilities of LLMs (Bareiss et al., 2022), visual programming frameworks (Surís et al., 2023; Gupta and Kembhavi, 2023; Subramanian et al., 2023) generate programs given an API and in-context examples and execute the programs with large pre-trained vision models. Visual programming has been applied to many different domains and applications including visual question answering, video question answering, text-to-image generation, and robotics (Cho et al., 2024; Lu et al., 2024; Min et al., 2024; Liang et al., 2023). Unlike previous visual programming works, we only use small models ($<1$ billion parameters) to generate programs.
Visual programming is different from both LLM tool use and LLM code generation since it requires both basic reasoning and knowledge skills as well as basic code generation. One of the challenging parts of visual programming comes from the execution and evaluation due to underlying models. As noted in other works Khan et al. (2024), incorrect programs can return (after being executed) the correct answer even if the program is incorrect.
@@
Compared to prompt-based methods, there are relatively few methods focused on improving program generation in visual programming or in the general field of tool-based LLMs through finetuning. One of the main challenges is the lack of program annotations for input/output pairs. Language modeling has a long history designing self-supervised tasks, especially in pre-training (Devlin et al., 2019; Raffel et al., 2020). Toolformer (Schick et al., 2024), uses a form of self-supervision, to create a training dataset to finetune an LLM on tool-use programs. For each question/answer pair in a pre-existing training dataset, an LLM is prompted to generate a corresponding program. If the program decreases the training loss (of the same LLM) then the program annotation is added to a new dataset. Then the same LLM is finetuned on the generated dataset. A similar approach is used in Chain-of-Thought (CoT) finetuning (Zhu et al., 2023), program correction methods such as VDebugger (Wu et al., 2024) and visual programming based finetuning methods (Khan et al., 2024) such as VisRep. LoRA (Hu et al., 2021) is frequently used when finetuning LLMs on the generated data.
In VisRep, Khan et al. (2024) use a self-training approach similar to toolformer for visual programming. Self-annotated programs are kept if the executed program returns the correct answer. While there are some similarities between VisRep and our work, the underlying setting is different. VisRep focuses on improving program generation on an existing LLM (7B) through self-training, while the goal of our work is to distill visual program generation from an LLM to a small model (<1B). Our work also does not require sampling across specific question types or correcting by hand different program annotations during training. We randomly sample the training set and do not write any program annotations or corrections.
@@
Prompt-based methods are more effective when in-context examples or demonstrations (Wei et al., 2022) are included. In visual programming (Surís et al., 2023; Gupta and Kembhavi, 2023), manually written in-context examples are used to adapt an API to a particular dataset. However, such an approach can be time consuming and not feasible for a large training set of input/output pairs. A common practice is to have an LLM self-annotate examples and keep the examples that produce the correct answer Zhang et al. (2023); Stanić et al. (2024); Tao et al. (2024); SU et al. (2023). Then during inference, augmented generation (RAG) Lewis et al. (2020)can be performed to select the most relevant ones. We follow such an approach, referred to as auto-context generation by  Tao et al. (2024) to generate in-context examples for the teacher model.
@@
3 Method
Preliminaries
Following the notation in ViperGPT Surís et al. (2023), the visual programming objective is to generate a program $z=\pi(q,p)$ with a program generator, $\pi$, input query $q$ and prompt $p$ such that when executed with execution engine $\phi$ and corresponding visual input $x$, $\phi(x,z)$ returns the correct answer. Prompt $p$ contains an API and dataset-specific in-context examples.
Given a teacher model, $\pi_{t}$, and a smaller student model ($<1$ billion parameters), $\pi_{s}$, our goal is to distill visual program generation from the teacher to the student for a specific dataset using a minimal number of answer annotations (question/answer pairs). We do not have any human provided annotations (question/program/answer triplets) but we do know the API used for the teacher model.
3.1 Teacher Generated Program Annotation
We use GPT-4o-Mini AI (2024) as our teacher model. Unlike traditional knowledge distillation, LLM based distillation requires the teacher (an LLM) to be given the appropriate context through in-context examples before starting annotation. Common practice is to have a human generate the examples. We follow a process known as Auto-Context Generation to automatically generate such examples given answer annotations only.
Figure 2: An example of our data augmentation approach. Both the new and old question have the same template so the template matcher output should predict the same template for both. The arguments for the new and old programs are different. But, in the arguments, (dog, sofa, brown) should be replaced with (bear, desk, green).
  1. 1.
  2. 2.
    Generated program is evaluated.
  3. 3.
For efficiency, the in-context examples are sampled based on similarity to the question (i.e. RAG) Lewis et al. (2020) once there are more than 50. We compute the cosine similarity with finetuned MP-Net-v2 Reimers and Gurevych (2019) between an input and all in-context examples and select the 50 highest scoring examples.
@@ 3.2 Data Augmentation
To understand the intuition behind our data augmentation method, consider the set of questions in  Table 2. All of these questions compare properties of two objects. The general structure of each program is the same with the only difference coming from the inputs to the functions. If we answer one of these questions correctly and know that the remaining questions have the same structure, then all of the remaining questions should have that same structure or should be consistent. An overview of our data augmentation method can be seen in Figure 1. We refer to this as ‘template-based augmentation.’
@@
We define a template as a specific ordering of functions, where a function is an API call to a visual model or python operation. Templates are argument independent. For example, if the program is
@@
Templates can be considered high-level plans used for plan-based distillation Sun et al. (2024). The main advantage to templates is that they can be extracted directly from the program. There is no need to generate extra output or perform multiple forward passes for a single question.
@@
Given the program annotations from the teacher model, we extract templates and corresponding arguments from each program similar to the abstraction method of CRAFT Yuan et al. (2024). Extracting is quite simple: replace specific variable names with generic ones and put in placeholders for each argument. The variable renaming and extraction can be done in seconds with abstract syntax trees and regular expression matching. The code can be found in Appendix E.
@@
We can use the decomposition of programs into templates and arguments to generate synthetic data similar to how masked language modeling is used in BERT Devlin et al. (2019). As can be seen in Table 2, for many questions and programs, the arguments appear directly in the question. Consider the example in Figure 2: “Are both the dog and the sofa brown?" The arguments are dog, sofa, and brown. Once we find similar words for each, we can simply replace them in the sentence. Since we already know the template and the arguments, we also have the program. For GQA (Hudson and Manning, 2019), the possible word replacements are from ViperGPT Surís et al. (2023) and for arguments that are either not in the question or are equal to the entire question, we use GPT to generate similar questions. For VQAv2 Goyal et al. (2017), we use BERT Devlin et al. (2019) when an argument is a single word and BART Lewis (2019) to replace phrases. If possible, we replace arguments with arguments of the same type e.g. an attribute is replaced by an attribute.
Example Template
var1 = image_patch.find(<arg_0>)
var2 = filter_img(<arg_1>)
var3 = var2.classify(<arg_2>)
var4 = image_patch.find(<arg_3>)
var5 = var4.classify(<arg_4>)
answer =bool_to_yesno(var3 ==absent$=$ = var5)
var1 = image_patch.find(<arg_0>)
var2 = var1.crop_position(<arg_1>)
var3 = var2.find(<arg_2>)
answer $=$ var3.classify(<arg_3>)
var1 = image_patch.find(<arg_0>)
var2 = image_patch.find(<arg_1>)
answer $=$ choose_relationship(var1,var2,<arg_2>)
During training, each argument in each question has a 50% chance of being replaced. If an argument is to be replaced, we then uniformly sample among the possible replacements. In the GQA dataset, for some arguments, like objects, the number of possible replacements is quite large (e.g. greater than 1500), while for arguments like directions such as left, right, etc. the number of replacements is small (e.g. fewer than 10). Some examples of categories and more details on word replacement can be found in Section F.
@@
Given the augmented dataset, we perform LoRA-based finetuning on the student model. Since the augmentation method produces both questions and corresponding programs, we use next-token prediction loss for training.

### v4

# Visual Program Distillation with Template-Based Augmentation

@@ Abstract
Adapting visual programming or prompting large language models (LLMs) to generate executable code for visual tasks like visual question answering (VQA) for specialized tasks or domains remains challenging due to high annotation and inference costs. We propose a low-cost visual program distillation method that can be used for models with at most 1 billion parameters and requires no human-generated program annotations. We achieve this through synthetic data augmentation based on decoupling programs into higher-level skills, called templates, and their corresponding arguments. Experimental results show that, with a relatively small amount of question/answer data, small language models can generate high-quality specialized visual programs with the added benefit of much faster inference.
Visual Program Distillation with Template-Based Augmentation
@@ 1 Introduction
Visual programming Gupta and Kembhavi (2023); Subramanian et al. (2023); Surís et al. (2023) refers to generating programs that invoke visual models to solve tasks such as answering questions about images, typically by prompting a very large language model (LLM) like GPT Achiam et al. (2023) or Llama Touvron et al. (2023).
Visual programming offers greater adaptability and customization for specialized applications such as a personal visual navigation assistant compared to a single vision-language model (VLM). In-context learning with proprietary or closed-source LLMs can generate correct visual programs for targeted applications but at the cost of long inference time and compute (see Figure 1). In addition, generating in-context examples (i.e. visual programs written by hand) requires a significant amount of human effort. Previous efforts Khan et al. (2024) have made some progress in adapting smaller, open-source LLMs for dataset-specific visual programs but still suffer from high training and data costs.
Figure 1: Accuracy vs. Throughput for Visual Program Generation on GQA Generalist LLMs (teacher models) offer high accuracy at the cost of low throughput and large model size (proportional to marker size). With our template-based augmentation method, specialized distilled student models achieve comparable performance on answer accuracy with a small percent of question/answer data ($\approx 0.1\%$) and no human program annotations.
Our goal is to create a specialized visual programming system with two key characteristics:
@@
    Small Program Generator Programs should be generated by models with $\leq 1$ billion parameters, enabling fast inference and use on a wide variety of hardware.
@@
    Limited Human Annotations Minimize the number of required annotations, especially human-generated programs, for ease of adaptation.
@@
Distill Step-by-Step Hsieh et al. (2023) Text Gen. LLM Distillation 220M-1B 12.5% Q/A, Reasoning
@@
Template-Based Aug. (Ours) Program Gen. LLM Distillation 500M-1B 0.1% ($\approx$1K) Visual Prog.
@@
Our key insight in achieving such capabilities lies in decoupling the skill or procedure from the question-specific concept. We call the higher-level skills templates and the concepts arguments. For example, the programs “Count the red chairs" and “Count the green bananas" have the same template,‘find(arg1), verify property(arg2), count’, but different arguments: ‘red’ (arg1) and ‘chairs’ (arg2) vs ‘green’ (arg1) and ‘bananas’ (arg2) respectively. This decomposition facilitates creating synthetic examples by replacing arguments in the question and program, e.g. ‘Count the red apples’ with corresponding argument substitution in the program. Given a small training dataset, we can use this process, which we refer to as template-based augmentation to increase concept diversity.
Combining template-based augmentation with recent advances in LLM distillation Hsieh et al. (2023), we propose a low-cost visual program distillation method. Our approach requires no human-generated example programs and uses only a small fraction of question/answer pairs (at most $0.1\%$ of the training dataset). A teacher model leverages auto-context generation, where each generated program that produces the correct answer is added to the set of in-context examples. The resulting annotated programs form a dataset of question/program pairs that are augmented via the template-based method described above and used to train a small language model (SLM) that efficiently generates visual programs.
Our method uses a low-cost teacher model (e.g. GPT-Mini-4o) and costs only around \$1 per dataset. In addition, the specialized distilled models achieve much faster inference speed (up to $\approx 30.8$x) than the teacher models.
To validate our approach, we evaluate on widely adopted visual question answering (VQA) datasets and compare with using human-generated annotations, few-shot prompting and non-augmented distillation. Most existing work evaluates only on whether the generated answer is correct. We find additional insights by also evaluating the correctness of the generated programs (as judged by a human) and student/teacher agreement, revealing substantial discrepancies among the different metrics.
We summarize our contributions as follows:
  • Introduce a template-based augmentation framework that distills large, generalist visual programming systems into small, specialized ones with minimal human effort. Template-based augmentation increases performance across all evaluated metrics.
  • Empirically show auto-context generation achieves the same or better performance as human provided annotations when prompting a teacher model.
  • Discuss and evaluate different metrics: answer accuracy, student/teacher agreement, and program accuracy for visual program evaluation. Surprisingly, we find the rate of error in answers is up to 3.8 times higher than that of the programs, suggesting that future visual programming work should focus on more effective APIs.
@@ 2 Related Works
A long line of work investigates generating and executing programs to perform visual tasks. Early approaches (Andreas et al., 2015, 2016; Hu et al., 2017; Johnson et al., 2017) generate programs and execute the programs with learned end-to-end neural modules. Based on the impressive code generation capabilities of LLMs (Bareiss et al., 2022), visual programming frameworks (Surís et al., 2023; Gupta and Kembhavi, 2023; Subramanian et al., 2023) generate programs given an API and in-context examples and execute the programs with large pre-trained vision models. Visual programming has been applied to many different domains and applications including visual question answering, video question answering, text-to-image generation, and robotics (Cho et al., 2024; Lu et al., 2024; Min et al., 2024; Liang et al., 2023). Unlike previous visual programming works, we only use small models ($\leq 1$ billion parameters) to generate programs.
One of the challenges that distinguishes visual programming from LLM tool use is the dependence of execution and evaluation on the output of underlying models. As noted in other works Khan et al. (2024), incorrect programs can return (after being executed) the correct answer even if the program is incorrect, which motivates our investigation into different evaluation metrics.
A concurrent work, Visual Unit Testing (ViUniT) Panagopoulou et al. (2025), generates visual unit tests for both visual program selection and self-training. ViUniT also evaluates the unit test quality and program correctness and reports similar differences (36%) between program and answer accuracy.
@@
Compared to prompt-based methods, there are relatively few methods focused on improving program generation in visual programming for specific tasks or in the general field of tool-based LLMs through finetuning. One of the main challenges is the lack of program annotations for input/output pairs. Language modeling has a long history designing self-supervised tasks, especially in pre-training (Devlin et al., 2019; Raffel et al., 2020). Toolformer (Schick et al., 2024), uses a form of self-supervision, to create a training dataset to finetune an LLM on tool-use programs. For each question/answer pair in a pre-existing training dataset, an LLM is prompted to generate a corresponding program. If the program decreases the training loss (of the same LLM), the program annotation is added to a new dataset. Then the same LLM is finetuned on the generated dataset. A similar approach is used in Chain-of-Thought (CoT) finetuning (Zhu et al., 2023), program correction methods for specific datasets such as VDebugger (Wu et al., 2024) and specialized visual programming methods (Khan et al., 2024), such as VisRep. LoRA (Hu et al., 2022) is frequently used when finetuning LLMs on the generated data.
In VisRep, Khan et al. (2024) use a self-training approach similar to toolformer to finetunue an LLM for dataset-specific visual programming. Self-annotated programs are kept if the executed program returns the correct answer. While there are some similarities between VisRep and our work, the underlying setting is different. VisRep focuses on improving program generation on an existing LLM (7B) through self-training for specialized tasks, while the goal of our work is to distill visual program generation from an LLM to a small model ($\leq 1$B) for specialized tasks. Our work also does not require sampling across specific question types or correcting by hand different program annotations during training. We randomly sample the training set and do not write any program annotations or corrections.
@@
Prompt-based methods are more effective when in-context examples or demonstrations (Wei et al., 2022) are included. In visual programming (Surís et al., 2023; Gupta and Kembhavi, 2023), manually written in-context examples are used to adapt an API to a particular dataset. However, such an approach can be time consuming and not feasible for a large training set of input/output pairs. A common practice is to have an LLM self-annotate examples and keep the examples that produce the correct answer Zhang et al. (2023); Stanić et al. (2024); Tao et al. (2024); SU et al. (2023). Then during inference, retrieval augmented generation (RAG) Lewis et al. (2020b) can be performed to select the most relevant ones. We follow such an approach, referred to as auto-context generation by  Tao et al. (2024) to generate in-context examples for the teacher model.
3 Method
Preliminaries
Following the notation in ViperGPT Surís et al. (2023), the visual programming objective is to generate a program $z=\pi(q,p)$ with a program generator $\pi$, input query $q$ and prompt $p$ such that when executed with an execution engine $\phi$ and corresponding visual input $x$, $\phi(x,z)$ returns the correct answer. Prompt $p$ contains an API and dataset-specific in-context examples.
Given a teacher model, $\pi_{t}$, and a smaller student model ($\leq 1$ billion parameters) $\pi_{s}$, our goal is to distill visual program generation from the teacher to the student for a specific dataset using a minimal number of answer annotations (question/answer pairs). We do not have any human provided program annotations (question/program/answer triplets), but we do know the API used for the teacher model.
3.1 Teacher Generated Program Annotation
We use GPT-4o-Mini Open AI (2024) as our teacher model since it is relatively cheap and has strong performance. Unlike traditional knowledge distillation, LLM based distillation requires the teacher (an LLM) to be given the appropriate context through in-context examples before starting annotation. Common practice is to have a human generate the examples. We follow a process known as auto-context generation Tao et al. (2024) to automatically generate such examples given answer annotations only.
  1. 1.
  2. 2.
    Generated program is evaluated.
  3. 3.
For efficiency, the in-context examples are sampled based on similarity to the question (i.e. RAG) Lewis et al. (2020b) once there are more than $x$ examples. We compute the cosine similarity with finetuned MP-Net-v2 Reimers and Gurevych (2019) between an input and all in-context examples and select the $x=50$ highest scoring examples.
@@ 3.2 Data Augmentation
To understand the intuition behind our data augmentation method, consider the set of questions in  Table 2. All of these questions compare properties of two objects. The general structure of each program is the same with the only difference coming from the inputs to the functions. If we answer one of these questions correctly and know that the remaining questions have the same structure, then all the remaining questions should also have that same structure or be consistent. An overview of our data augmentation method can be seen in Figure 2. We refer to this as "template-based augmentation."
Example Template
var1 = image_patch.find(<arg_0>)
var2 = filter_img(<arg_1>)
var3 = var2.classify(<arg_2>)
var4 = image_patch.find(<arg_3>)
var5 = var4.classify(<arg_4>)
answer =bool_to_yesno(var3 ==== var5)
var1 = image_patch.find(<arg_0>)
var2 = var1.crop_position(<arg_1>)
var3 = var2.find(<arg_2>)
answer == var3.classify(<arg_3>)
var1 = image_patch.find(<arg_0>)
var2 = image_patch.find(<arg_1>)
answer == choose_relationship(var1,var2,<arg_2>)
@@
We define a template as a specific ordering of functions, where a function is an API call to a visual model or python operation. Templates encode program structure independently of argument values. For example, if the program is
@@
Templates can be considered high-level plans used for plan-based distillation Sun et al. (2024). The main advantage of templates is that they can be extracted directly from the program. There is no need to generate extra output or perform multiple forward passes for a single question.
Figure 3: An example of our data augmentation approach. Both the new and old question have the same template, so the template matcher output should predict the same template for both. The arguments for the new and old programs are different. But, in the arguments, (dog, sofa, brown) should be replaced with (bear, desk, green).
@@
Given the program annotations from the teacher model, we extract templates and corresponding arguments from each program similar to the abstraction method of CRAFT Yuan et al. (2024). Extracting is quite simple: replace specific variable names with generic ones and put in placeholders for each argument. The variable renaming and extraction can be done in seconds with abstract syntax trees and regular expression matching. Note that while the template extraction algorithm is deterministic, the actual templates are determined solely by the annotations from the teacher model. The code can be found in Appendix F.
@@
We can use the decomposition of programs into templates and arguments to generate synthetic data similar to how masked language modeling is used in BERT Devlin et al. (2019). As can be seen in Table 2, for many questions and programs, the arguments appear directly in the question. Consider the example in Figure 3: "Are both the dog and the sofa brown?" The arguments are dog, sofa, and brown. Once we find similar words for each, we can simply replace them in the sentence. At this point, we also have the program, since the template and arguments are known.
There are two different methods for generating word replacements. For the GQA (Hudson and Manning, 2019) dataset, the possible word replacements come from ViperGPT Surís et al. (2023) except for two special circumstances. The first is when arguments are not in the question. For example, the question "What is that made from?" could have a program where the arguments are "object" (with the find function) and "material" (with the classify function). The second is when the program consists of only calling a VLM and the input is the entire question. For the special cases, we use GPT Open AI (2024) to generate similar questions. For VQAv2 Goyal et al. (2017), we use BERT Devlin et al. (2019) when an argument is a single word and BART Lewis et al. (2020a) to replace phrases. If possible, we replace arguments with arguments of the same type, e.g. an attribute is replaced by an attribute.
During training, each argument in each question has a 50% chance of being replaced. If an argument is to be replaced, we then uniformly sample among the possible replacements. In the GQA dataset, for some arguments, like objects, the number of possible replacements is quite large (e.g. greater than 1500), while for arguments like directions such as left, right, etc. the number of replacements is small (e.g. fewer than 10). Some examples of categories and more details on word replacement can be found in Appendix G.
@@
Given the augmented dataset, we perform LoRA-based finetuning on the student model. The input to the student model is the question and the output is the visual program. Since the augmentation method produces both questions and corresponding programs, we use a next-token prediction loss for training.